Arnold had $1.70 in dimes and quarters. HE had 3 more dimes than quarter. How many of each coin did he have?
Answer by tanimachatterjee(60)

suppose Arnold had x dimes and y quarters
value of x dimes = 10x cents
y quarters =25y cents
tolal value =10x+25y cents
by the problem 10x+ 25y=170..............(1)
and x=y+3....(2)
replacing the value of x in (1) we have 10(y+3) +25y=170
or 10y +30 +25y=170
or 35y= 170-30=140
y= 4
x=y+3
= 4+3
=7
Arnold had 7 dimes and 4 quarters